OpenStudy (anonymous):

Find the horizontal asymptote of y= 1/x-3 + 2. I just don't get the whole asymptote thing.

OpenStudy (tkhunny):

You have written: \(y = \dfrac{1}{x} - 3 + 2\) I'm guessing this is not your intent. Give it another go and remember your Order of Operations.

OpenStudy (anonymous):

Sorry I meant \[y= \frac{ 1 }{ x-3 } + 2\]

OpenStudy (tkhunny):

Consider the degree of the numerator and denominator separately. What say you?

OpenStudy (anonymous):

Num deg = 0, denom. deg = 1

OpenStudy (anonymous):

So, since the numerator is a lower degree than the denominator, then there is a horizontal asymptote y= 0.

OpenStudy (anonymous):

So do I plug that in?

OpenStudy (anonymous):

\[0 =\frac{ 1 }{x-3}+ 2\]

OpenStudy (anonymous):

????????

OpenStudy (tkhunny):

You see a horizontal asymptote. Why are you substituting anything? \(y = \dfrac{1}{x-3}\) has a horizontal asymptote at y = 0 Move it up with a translation of 2. \(y = \dfrac{1}{x-3} + 2\) has a horizontal asymptote at y = 2

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOOOOOOH!

OpenStudy (anonymous):

THANK YOU!

OpenStudy (tkhunny):

You can alsow rewrite, by adding the fractions. \(y = \dfrac{1}{x-3} + 2 = \dfrac{2x-5}{x-3}\). The asymptote at y = 2 is more obvious in this form.